how far would a car go in 8 hours if it runs 4.5 km/hr.?

Answer by lmeeks54(111)

This is pretty straightforward and requires understanding the relationships between distance (d), time (t), and rate (r).
...
The basic formula for distance, rate, and time is:
...
d = r * t
...
Given:
t = 8 hrs
r = 4.5 km/hr
...
Solve for d:
...
d = 4.5 km/hr * 8 hrs
d = 36 km (the hours cancel out)
...
Note: for other problems in this same series asking you to solve for r or t, rewrite the basic equation and solve for the missing quantity:
...
d = r * t
r = d / t
t = d / r
...
Cheers,
Lee